OpenStudy (anonymous):

Solve the equation 6D-3D+2D-6=4D A.) 2 B.)6 C.) -8 D.) 2/3

OpenStudy (anonymous):

Hey Drea! hope this helps 6d - 3d + 2d - 6 = 4d 3d + 2d - 6 = 4d 5d - 6 = 4d 5d - 4d - 6 = 0 d - 6 = 0 d = 6